# Spectral theory of the Neumann-Poincaré operator associated with multi-layer structures and analysis of plasmon mode splitting

### v2

@@ Abstract
In this paper, we develop a general mathematical framework for analyzing electostatics within multi-layer metamaterial structures. The multi-layer structure can be designed by nesting complementary negative and regular materials together, and it can be easily achieved by truncating bulk metallic material in a specific configuration. Using layer potentials and symmetrization techniques, we establish the perturbation formula in terms of Neumann-Poincaré (NP) operator for general multi-layered medium, and obtain the spectral properties of the NP operator, which demonstrates that the number of plasmon modes increases with the number of layers. Based on Fourier series, we present an exact matrix representation of the NP operator in an apparently unsymmetrical structure, exemplified by multi-layer confocal ellipses. By highly intricate and delicate analysis, we establish a handy algebraic framework for studying the splitting of the plasmon modes within multi-layer structures. Moreover, the asymptotic profiles of the plasmon modes are also obtained. This framework helps reveal the effects of material truncation and rotational symmetry breaking on the splitting of the plasmon modes, thereby inducing desired resonances and enabling the realization of customized applications.
Key words: Neumann-Poincaré operator; multi-layer structures; plasmon mode splitting; characteristic polynomial
@@ 1 Introduction
We are concerned with the mathematical investigation of electostatics within multi-layer structure and the peculiar plasmon resonance phenomena it induces. Plasmon resonance refers to the resonant oscillation of conduction electrons at the interface between negative and positive permittivity material caused by the background field. Negative materials can be created using high-contrast particles compared to the surrounding medium through homogenization theory in specific configurations [16, 9]. Another commonly utilized category of negative materials includes noble metallic materials [31, 38]. Negative materials have been proposed for groundbreaking applications such as super-resolution imaging and invisibility cloaking. Multi-layer structures are frequently employed as the foundational components in constructing metamaterial devices for these cutting-edge applications [41]. It is important to note that the plasmon resonances induced by these negative structures serve as the fundamental basis for these applications.
@@
In [5, 8], the authors studied the surface plasmon resonance of metallic nanoparticles in electromagnetic wave using the Drude model, with applications in wave imaging. These nanoparticles can be viewed as single-layer structures. In [39, 4, 3, 15], the core-shell structures were proposed for achieving invisibility cloaking through the phenomenon known as anomalous localized resonance. In this configuration, the core contains a regular material while the shell contains a negative material, both being uniform. By precisely aligning their material properties according to specific geometric arrangements, cloaking due to anomalous localized resonance can be induced by some appropriate incident field. The core-shell device can be regarded as a two-layer structure. The aforementioned studies have also been extended in [33, 34] for macro-scale scenarios, in [43, 42, 22, 23] for anisotropic structures (nanorod, slender-body), and in [18] for two-layer confocal ellipses. Similarly, one-layer or two-layer metamaterial structures have also been extensively used and theoretically investigated in linear elasticity, see e.g. [20, 35, 36, 11, 10, 21] and the references cited therein. Multiband operation and cloaking of larger objects have been achieved using multi-layer plasmonic covers, providing more degrees of freedom to suppress a greater number of scattering orders [1, 17]. Consequently, three-layer or four-layer structures, known as nanomatryushkas (a matryushka is a nested set of dolls found in Russian folk art), have been widely employed and theoretically investigated using the hybridization model [41, 25, 40, 32] and the references cited therein. Notably, the plasmon resonance of multi-layer concentric radial structures has been investigated in [24, 30], where plasmon modes interact and hybridize, resulting in localized surface plasmons excited at each surface. This phenomenon explains the physical reason why the number of plasmon modes splitting (PMS) corresponds to the number of layers in the structure. This would enable an even greater degree of versatility in designing the cloak, accommodating multiple operating frequencies and potentially wider frequency bandwidths.
@@
Despite the important work mentioned above, there are still many outstanding questions regarding PMS within multi-layer structures waiting to be answered. One such question deals with PMS on multi-layer structures of general shape: most of the aforementioned works focus on radial structures. As the first step to addressing this challenging question, we consider in this paper the structure of multi-layer confocal ellipses. The number of layers can be arbitrary and the material parameters in each layer may be different, though uniform. This can encompass many of the existing studies mentioned above, where the number of layers is one or two. However, these results cannot be easily generalized to arbitrary multi-layer structures, which poses a significant challenge. To the authors’ best knowledge, the physical reason for the PMS within multi-layer confocal ellipses has not been adequately discussed in the literature and neither has the equivalence of the multi-layer circular and elliptical structures. We prove that like the radial case there is an exact perturbed field in terms of the generalized polarization matrix (GPM) which contains the material information, together with the geometric information of multi-layer structures. The explicit formula for the determinant of the GPM, which is equivalent to the characteristic polynomial of matrix type Neumann-Poincaré (NP) operator in multi-layer structures, is a stumbling block in the analysis of PMS.
@@
To prove the main result of this paper, we first derive the matrix type NP operator $\mathbb{K}_{N}^{*}$ (see (2.13)) in the $N$-layer structure with $C^{1,\eta}\,(0<\eta<1)$ smooth interface using layer potentials. Secondly, by delicate analysis, we develop a spectral theory (see Proposition 3.1) of NP operator, especially the Calderón-type identity in multi-layer structures, which serves as a highly nontrivial extension of the corresponding study about two-layer structures in [4]. Based on computations of the NP operator on single-layer ellipses in [7], we obtain two exact matrix representation of $\mathbb{K}_{N}^{*}$: even NP matrix $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$ and odd NP matrix $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$ (see Corollary 5.1), which results from a breaking of rotational symmetry; however the modes specifically split into even and odd parities because the multi-layer confocal ellipses still possesse two mirror planes. After struggling against complicated structure of the matrics $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$ and $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$, we come up with a conjecture of the exact formulation of characteristic polynomial. The conjecture is then verified by backward induction. By using the spectral theory of NP operator, we conclude that the even NP matrix $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$ and odd NP matrix $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$ have $N$ real eigenvalues that lie in the interval $[-1/2,1/2]$, respectively. Each eigenvalue is associated with one type of plasmon modes, which means that truncating the material will result in plasmon mode splitting. Furthermore, the asymptotic profiles of the plasmon modes in multi-layer confocal ellipses are obtained as the elliptic radii approach infinity (i.e., multi-layer concentric discs) or zero (i.e., multi-layer thin strips). This analysis establishes an algebraic framework that offers a convenient and practical approach for studying plasmon resonances associated with multi-layer metamaterial structures. In practical applications, the multi-layer structure can serve as a fundamental building block for various material devices. Our algebraic framework will provide a powerful and general design principle that can be applied to select appropriate material parameters, both qualitatively and quantitatively, guiding the design of metallic nanostructures and predicting their resonant properties. We shall investigate along this direction in our forthcoming work.
@@
The remainder of this paper is organized as follows. In Section 2, we first present the electrostatic scattering problem with multi-layer structures. Second, we transform the problem into an integral system using layer potentials and derive the form of the NP operator in multi-layer structures. In Section 3, we develop a spectral theory for NP operator. In Section 4, we give the representation of the perturbed field in terms of the GPM. In Section 5, we establish an algebraic framework for plasmon resonances and give the main results on the explicit formula of the characteristic polynomial and the estimation of the roots. Section 6 is devoted to the proofs of the main results in Section 5. In Section 7, numerical examples are presented in finding all the plasmon resonance modes in a fixed multi-layer structure and the surface localization and geometrization of plasmon resonant modes is simulated. Some conclusions are made in Section 8.
2 Electrostatics with multi-layer structures
2.1 Multi-layer structures
@@
In this subsection, we first introduce the material parameters and geometric configurations for subsequent analysis. The definition below encapsulates the concept of general multi-layer structures.
@@
Let $A$ be the inclusion with a $C^{1,\eta}\,(0<\eta<1)$ smooth boundary $\Gamma_{1}:=\partial A$ and $A_{0}=\mathbb{R}^{d}\setminus\overline{A}$. We say that $A$ is considered to have a partition made up of a multi-layer structure, if the interior of $A$ be divided by means of closed and nonintersecting $C^{1,\eta}$ surfaces $\Gamma_{k}$ $(k=2,3,...,N)$ into subsets (layers) $A_{k}$ $(k=1,2,...,N)$. Each $\Gamma_{k-1}$ surrounds $\Gamma_{k}$ and $A_{k}\subset\subset(A_{k-1}\cup A_{k})$, $(k=2,3,\ldots,N)$. The region $A_{k}$ $(k=0,1,2,\ldots,N)$ stands for homogeneous media.
@@
$\sigma(x)=\sigma_{c}(x)\chi(A)+\sigma_{0}\chi(\mathbb{R}^{d}\backslash\overline{A}),$ (2.1)
where $\chi$ stands for the characteristic function of a domain; $\sigma_{0}$ is set to be a positive constant and hence it defines a regular material in the background space $\mathbb{R}^{d}\backslash\overline{A}$. We set $\sigma_{c}$ to be of the following form
@@
In summary, we consider a rather general multi-layer structure in which the number of layers can be arbitrarily given and each layer can consist of either regular or negative materials. Moreover, the material parameters in each layer can be different from one another.
@@ 2.2 Electrostatics scattering and integral reformulation
where $d\geqslant 2$, the background electrical potential ${H}$ is a harmonic function in $\mathbb{R}^{d}$, and $u$ represents the total electric potential. It is nature that the solution $u$ to (2.3) satisfies the transmission conditions
@@
Define $\mathcal{H}^{s}=\prod_{k=1}^{N}\otimes H^{s}(\Gamma_{k})$ and $\mathcal{H}^{s}_{0}=\prod_{k=1}^{N}\otimes H^{s}_{0}(\Gamma_{k})$. With the above preparations, one has the following integral representation of the conductivity transmission problem (2.1)–(2.4) with density function $\phi_{k}\in H^{-\frac{1}{2}}_{0}(\Gamma_{k})$
@@
$\begin{split}\begin{bmatrix}\lambda_{1}I-\mathcal{K}_{\Gamma_{1}}^{*}&-\nu_{1}\cdot\nabla\mathcal{S}_{\Gamma_{2}}&\cdots&-\nu_{1}\cdot\nabla\mathcal{S}_{\Gamma_{N}}\\ -\nu_{2}\cdot\nabla\mathcal{S}_{\Gamma_{1}}&\lambda_{2}I-\mathcal{K}_{\Gamma_{2}}^{*}&\cdots&-\nu_{2}\cdot\nabla\mathcal{S}_{\Gamma_{N}}\\ \vdots&\vdots&\ddots&\vdots\\ -\nu_{N}\cdot\nabla\mathcal{S}_{\Gamma_{1}}&-\nu_{N}\cdot\nabla\mathcal{S}_{\Gamma_{2}}&\cdots&\lambda_{N}I-\mathcal{K}_{\Gamma_{N}}^{*}\end{bmatrix}\begin{bmatrix}\phi_{1}\\ \phi_{2}\\ \vdots\\ \phi_{N}\end{bmatrix}=\begin{bmatrix}\nu_{1}\cdot\nabla H\\ \nu_{2}\cdot\nabla H\\ \vdots\\ \nu_{N}\cdot\nabla H\end{bmatrix},\end{split}$ (2.8)
@@
The problem (2.7) is solvable for $|\lambda_{k}|>\frac{1}{2}$ as shown in [29, Theorem 2.1], with the positivity assumption on the conductivities $\sigma_{k}$, $k=0,1,\ldots,N$. However, in plasmon modes, it contains negative values in the multi-layer structure, that is $\sigma_{k}<0$ holds for some $k=1,2,\ldots,N.$ This causes a serious difficulty in dealing with (2.8).
It is known that plasmon resonance is usually associated with some eigenvalue problem generated by the PDE system. We shall also explore the related eigenvalue problem for multi-layer structure. To simplify the analysis, we suppose that
@@
$\begin{split}\begin{bmatrix}\lambda I-\mathcal{K}_{\Gamma_{1}}^{*}&-\nu_{1}\cdot\nabla\mathcal{S}_{\Gamma_{2}}&\cdots&-\nu_{1}\cdot\nabla\mathcal{S}_{\Gamma_{N}}\\ \nu_{2}\cdot\nabla\mathcal{S}_{\Gamma_{1}}&\lambda I+\mathcal{K}_{\Gamma_{2}}^{*}&\cdots&\nu_{2}\cdot\nabla\mathcal{S}_{\Gamma_{N}}\\ \vdots&\vdots&\ddots&\vdots\\ (-1)^{N}\nu_{N}\cdot\nabla\mathcal{S}_{\Gamma_{1}}&(-1)^{N}\nu_{N}\cdot\nabla\mathcal{S}_{\Gamma_{2}}&\cdots&\lambda I+(-1)^{N}\mathcal{K}_{\Gamma_{N}}^{*}\end{bmatrix}\begin{bmatrix}\phi_{1}\\ \phi_{2}\\ \vdots\\ \phi_{N}\end{bmatrix}=\begin{bmatrix}\nu_{1}\cdot\nabla H\\ -\nu_{2}\cdot\nabla H\\ \vdots\\ (-1)^{N-1}\nu_{N}\cdot\nabla H\end{bmatrix},\end{split}$ (2.12)
Let $\mathbb{K}_{N}^{*}$ be an $N$-by-$N$ matrix type NP operator on $\mathcal{H}^{-1/2}_{0}$ defined by
$\begin{split}\mathbb{K}_{N}^{*}:=\begin{bmatrix}-\mathcal{K}_{\Gamma_{1}}^{*}&-\nu_{1}\cdot\nabla\mathcal{S}_{\Gamma_{2}}&\cdots&-\nu_{1}\cdot\nabla\mathcal{S}_{\Gamma_{N}}\\ \nu_{2}\cdot\nabla\mathcal{S}_{\Gamma_{1}}&\mathcal{K}_{\Gamma_{2}}^{*}&\cdots&\nu_{2}\cdot\nabla\mathcal{S}_{\Gamma_{N}}\\ \vdots&\vdots&\ddots&\vdots\\ (-1)^{N}\nu_{N}\cdot\nabla\mathcal{S}_{\Gamma_{1}}&(-1)^{N}\nu_{N}\cdot\nabla\mathcal{S}_{\Gamma_{2}}&\cdots&(-1)^{N}\mathcal{K}_{\Gamma_{N}}^{*}\end{bmatrix},\end{split}$ (2.13)

### v3

# Spectral theory of the Neumann-Poincaré operator on multi-layered structures and analysis of plasmon mode splitting

@@ Abstract
In this paper, we develop a general mathematical framework for analyzing electostatics within multi-layered metamaterial structures. The multi-layered structure can be designed by nesting complementary negative and regular materials together, and it can be easily achieved by truncating bulk metallic material in a specific configuration. Using layer potentials and symmetrization techniques, we establish the perturbation formula in terms of Neumann-Poincaré (NP) operator for general multi-layered medium, and obtain the spectral properties of the NP operator, which demonstrates that the number of plasmon modes increases with the number of layers. Based on Fourier series, we present an exact matrix representation of the NP operator in an apparently unsymmetrical structure, exemplified by multi-layered confocal ellipses. By highly intricate and delicate analysis, we establish a handy algebraic framework for studying the splitting of the plasmon modes within multi-layered structures. Moreover, the asymptotic profiles of the plasmon modes are also obtained. This framework helps reveal the effects of material truncation and rotational symmetry breaking on the splitting of the plasmon modes, thereby inducing desired resonances and enabling the realization of customized applications.
Key words: Neumann-Poincaré operator; multi-layered structures; plasmon mode splitting; characteristic polynomial
@@ 1 Introduction
We are concerned with the mathematical investigation of electostatics within multi-layered structure and the peculiar plasmon resonance phenomena it induces. Plasmon resonance refers to the resonant oscillation of conduction electrons at the interface between negative and positive permittivity material caused by the background field. Negative materials can be created using high-contrast particles compared to the surrounding medium through homogenization theory in specific configurations [16, 9]. Another commonly utilized category of negative materials includes noble metallic materials [31, 38, 47]. Negative materials have been proposed for groundbreaking applications such as super-resolution imaging and invisibility cloaking. Multi-layered structures are frequently employed as the foundational components in constructing metamaterial devices for these cutting-edge applications [41, 46]. It is important to note that the plasmon resonances induced by these negative structures serve as the fundamental basis for these applications.
@@
In [5, 8], the authors studied the surface plasmon resonance of metallic nanoparticles in electromagnetic wave using the Drude model, with applications in wave imaging. These nanoparticles can be viewed as single-layer structures. In [39, 4, 3, 15], the core-shell structures were proposed for achieving invisibility cloaking through the phenomenon known as anomalous localized resonance. In this configuration, the core contains a regular material while the shell contains a negative material, both being uniform. By precisely aligning their material properties according to specific geometric arrangements, cloaking due to anomalous localized resonance can be induced by some appropriate incident field. The core-shell device can be regarded as a two-layer structure. The aforementioned studies have also been extended in [33, 34] for macro-scale scenarios, in [43, 42, 22, 23] for anisotropic structures (nanorod, slender-body), and in [18] for two-layer confocal ellipses. Similarly, one-layer or two-layer metamaterial structures have also been extensively used and theoretically investigated in linear elasticity, see e.g. [20, 35, 36, 11, 10, 21] and the references cited therein. Multiband operation and cloaking of larger objects have been achieved using multi-layered plasmonic covers, providing more degrees of freedom to suppress a greater number of scattering orders [1, 17]. Consequently, three-layer or four-layer structures, known as nanomatryushkas (a matryushka is a nested set of dolls found in Russian folk art), have been widely employed and theoretically investigated using the hybridization model [41, 25, 40, 32] and the references cited therein. Notably, the plasmon resonance of multi-layered concentric radial structures has been investigated in [24, 30], where plasmon modes interact and hybridize, resulting in localized surface plasmons excited at each surface. This phenomenon explains the physical reason why the number of plasmon modes splitting (PMS) corresponds to the number of layers in the structure. This would enable an even greater degree of versatility in designing the cloak, accommodating multiple operating frequencies and potentially wider frequency bandwidths.
@@
Despite the important work mentioned above, there are still many outstanding questions regarding PMS within multi-layered structures waiting to be answered. One such question deals with PMS on multi-layered structures of general shape: most of the aforementioned works focus on radial structures. As the first step to addressing this challenging question, we consider in this paper the structure of multi-layered confocal ellipses. The number of layers can be arbitrary and the material parameters in each layer may be different, though uniform. This can encompass many of the existing studies mentioned above, where the number of layers is one or two. However, these results cannot be easily generalized to arbitrary multi-layered structures, which poses a significant challenge. To the authors’ best knowledge, the physical reason for the PMS within multi-layered confocal ellipses has not been adequately discussed in the literature and neither has the equivalence of the multi-layered circular and elliptical structures. We prove that like the radial case there is an exact perturbed field in terms of the generalized polarization matrix (GPM) which contains the material information, together with the geometric information of multi-layered structures. The explicit formula for the determinant of the GPM, which is equivalent to the characteristic polynomial of matrix type Neumann-Poincaré (NP) operator in multi-layered structures, is a stumbling block in the analysis of PMS.
@@
To prove the main result of this paper, we first derive the matrix type NP operator $\mathbb{K}_{N}^{*}$ (see (2.13)) in the $N$-layer structure with $C^{1,\eta}\,(0<\eta<1)$ smooth interface using layer potentials. Secondly, by delicate analysis, we develop a spectral theory (see Proposition 3.1) of NP operator, especially the Calderón-type identity in multi-layered structures, which serves as a highly nontrivial extension of the corresponding study about two-layer structures in [4]. Based on computations of the NP operator on single-layer ellipses in [7], we obtain two exact matrix representation of $\mathbb{K}_{N}^{*}$: even NP matrix $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$ and odd NP matrix $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$ (see Corollary 5.1), which results from a breaking of rotational symmetry; however the modes specifically split into even and odd parities because the multi-layered confocal ellipses still possesse two mirror planes. After struggling against complicated structure of the matrics $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$ and $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$, we come up with a conjecture of the exact formulation of characteristic polynomial. The conjecture is then verified by backward induction. By using the spectral theory of NP operator, we conclude that the even NP matrix $\left(\mathbf{K}_{N,c}^{(n)}\right)^{T}$ and odd NP matrix $\left(\mathbf{K}_{N,s}^{(n)}\right)^{T}$ have $N$ real eigenvalues that lie in the interval [1/2,1/2][-1/2,1/2], respectively. Each eigenvalue is associated with one type of plasmon modes, which means that truncating the material will result in plasmon mode splitting. Furthermore, the asymptotic profiles of the plasmon modes in multi-layered confocal ellipses are obtained as the elliptic radii approach infinity (i.e., multi-layered concentric discs) or zero (i.e., multi-layered thin strips). This analysis establishes an algebraic framework that offers a convenient and practical approach for studying plasmon resonances associated with multi-layered metamaterial structures. In practical applications, the multi-layered structure can serve as a fundamental building block for various material devices. Our algebraic framework will provide a powerful and general design principle that can be applied to select appropriate material parameters, both qualitatively and quantitatively, guiding the design of metallic nanostructures and predicting their resonant properties. We shall investigate along this direction in our forthcoming work.
@@
The remainder of this paper is organized as follows. In Section 2, we first present the electrostatic scattering problem with multi-layered structures. Second, we transform the problem into an integral system using layer potentials and derive the form of the NP operator in multi-layered structures. In Section 3, we develop a spectral theory for NP operator. Section 4 presents the representation of the perturbed field in terms of the GPM. In Section 5, we establish an algebraic framework for plasmon resonances and give the main results on the explicit formula of the characteristic polynomial and the estimation of the roots. Section 6 is devoted to the proofs of the main results in Section 5. In Section 7, numerical examples are presented to corroborate our theoretical results. Some conclusions are made in Section 8.
2 Electrostatics with multi-layered structures
2.1 Multi-layered structures
@@
In this subsection, we first introduce the material parameters and geometric configurations for subsequent analysis. The definition below encapsulates the concept of general multi-layered structures.
@@
Let $A$ be the inclusion with a $C^{1,\eta}\,(0<\eta<1)$ smooth boundary $\Gamma_{1}:=\partial A$ and $A_{0}=\mathbb{R}^{d}\setminus\overline{A}$. We say that $A$ is considered to have a partition made up of a multi-layered structure, if the interior of $A$ be divided by means of closed and nonintersecting $C^{1,\eta}$ surfaces $\Gamma_{k}$ (k=2,3,,N)(k=2,3,...,N) into subsets (layers) $A_{k}$ (k=1,2,,N)(k=1,2,...,N). Each $\Gamma_{k-1}$ surrounds $\Gamma_{k}$ $(k=2,3,\ldots,N)$. The region $A_{k}$ $(k=0,1,2,\ldots,N)$ stands for homogeneous media.
@@
[…]
where $\chi$ stands for the characteristic function of a domain; $\sigma_{0}$ is set to be a positive constant and hence it defines a regular material in the background space $\A¯\mathbb{R}^{d}\backslash\overline${A}. We set $\sigma_{c}$ to be of the following form
@@
In summary, we consider a rather general multi-layered structure in which the number of layers can be arbitrarily given and each layer can consist of either regular or negative materials. Moreover, the material parameters in each layer can be different from one another.
@@ 2.2 Electrostatics scattering and integral reformulation
where $d\geqslant 2$, the background electrical potential H{H} is a harmonic function in $\mathbb{R}^{d}$, and $u$ represents the total electric potential. It is nature that the solution $u$ to (2.3) satisfies the transmission conditions
@@
Define $\mathcal{H}^{s}=\prod_{k=1}^{N}H^{s}(\Gamma_{k})$ and $\mathcal{H}^{s}_{0}=\prod_{k=1}^{N}H^{s}_{0}(\Gamma_{k})$. With the above preparations, one has the following integral representation of the conductivity transmission problem (2.1)–(2.4) with density function $\phi_{k}\in H^{-\frac{1}{2}}_{0}(\Gamma_{k})$
@@
[…]
@@
The problem (2.7) is solvable for $|\lambda_{k}|>\frac{1}{2}$ as shown in [29, Theorem 2.1], with the positivity assumption on the conductivities $\sigma_{k}$, $k=0,1,\ldots,N$. However, in plasmon modes, it contains negative values in the multi-layered structure, that is $\sigma_{k}<0$ holds for some $k=1,2,\ldots,N.$ This causes a serious difficulty in dealing with (2.8).
It is known that plasmon resonance is usually associated with some eigenvalue problem generated by the PDE system. We shall also explore the related eigenvalue problem for multi-layered structure. To simplify the analysis, we suppose that
@@
[…]
Let $\mathbb{K}_{N}^{*}$ be an NN-by-NN matrix type NP operator on $\mathcal{H}^{-1/2}_{0}$ defined by
[…]